What is the z-score multiplier to make a 2-sided 95% confidence interval? (this is the value you look up in the standard normal table - enter your answer with 2 decimal places)?

What is the z-score multiplier used to make a 1-sided 95% confidence interval? (this is the value you look up in the standard normal table - enter your answer with 2 decimal places)?
Answer #1

\alpha = 1 - 0.95 = 0.05

For Two tailed test,

Z\alpha/2 = Z0.05/2 = Z0.025 = 1.96

\alpha = 1 - 0.95 = 0.05

For one tailed test,

Z\alpha = Z0.05 = 1.64